OpenStudy (anonymous):

If one side of a square is x-2, what is the value of x if the perimeter is 6x+7? Help

OpenStudy (anonymous):

Here you are given a two equations with which you can create a system to solve. You only need to translate the second one. How can you write an equations which connects the perimeter with one of the sides of the square ?

OpenStudy (anonymous):

4 times the x-2?

OpenStudy (anonymous):

4(x-2)=6x+7?

OpenStudy (anonymous):

I think im wrong

OpenStudy (anonymous):

Yes, thats it !

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Yes, now you have to solve the system.

OpenStudy (anonymous):

X=-15/2?

OpenStudy (anonymous):

Can you show the full way of going there ?

OpenStudy (anonymous):

Is it wrong?

OpenStudy (anonymous):

Yes. The answer can't be a negative number because if x is negative, then x-2 will also be negative and therefore the side of the square will be negative which can't actually happen

OpenStudy (anonymous):

Can you help me find the answer? I don't get this :(

OpenStudy (anonymous):

You need to solve 4( x-2 ) = 6x + 7 right ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Lets see..

OpenStudy (anonymous):

4( x-2 ) = 6x + 7 <=> 4x - 8 = 6x + 7 <=> 2x = - 15 <=> x = - 15 / 2

OpenStudy (anonymous):

So, there is something wrong about the problem :)

OpenStudy (anonymous):

It's on the radical review, I think it has something to do with radicals

OpenStudy (anonymous):

Dude, if the one side of the square is negative then there is no square. You can see yourself that if we assume that our square can have a side of - 15 / 2 - 2 = - 19 / 2 then its perimeter is 4 * ( - 19 / 2 ) = - 19 * 2 = - 38 But - 38 can't be the value of a perimeter although - 38 = 6 * ( - 15 / 2 ) + 7